# LIME: Less Is More for MLLM Evaluation

### v2

@@ 5.2 Correlation Analysis
Figure 4 illustrates the correlation between the various sub-tasks in LIME and WildVision Bench. Most tasks in LIME exhibit a strong positive correlation with WildVision Bench. There are six subtasks that have correlation scores exceeding 80%. Additionally, the overall score of LIME correlates at 91% with WV-Elo, which is higher than any individual sub-task and the original bench’s correlates, demonstrating that the overall score of LIME provides a more comprehensive reflection of MLLMs capabilities.
@@
As an earliest foundational problem, the captioning task has been extensively studied and MLLMs have an exceptional ability on this task. For instance, earlier models like InstructBlip performed exceptionally well on captioning tasks, and there is a broad presence of training data for image captioning in MLLMs’ training processes. However, the captioning task shows a negative correlation with all other sub-tasks. This indicates that previous metrics (e.g., BLEU, CIDEr) only focus on the overlap between the model-generated responses and the ground truth, but do not consider that MLLMs might generate content that is semantically close to the ground truth (i.e., the model-generated response may be semantically similar to the ground truth but expressed differently, or the model may generate more detailed content about the image). Consequently, we exclude it from the overall score calculation.
@@
On the one hand, the relevance of TextVQA, InfographicVQA and OCRBench are relatively high. As shown in the Tab 4, the correlation of those tasks all surpass 85%, and these two VQA tasks require MLLMs to understand the fine-grained content in images to answer question. It demonstrates that OCR tasks also rely on the ability of MLLMs to perceive fine-grained objective facts in images. On the other hand, POPE, ChartQA and InfographicVQA all need the reasoning ability using extra commonsense knowledge. The correlation scores of these tasks are all above 70%, and the POPE requires model to use the extra knowledge to solve the hallucination of MLLMs. We assume that ChartQA and infoVQA may also necessitate the use of additional common knowledge by the models to solve problems.
@@ 5.3 Effectiveness of LIME
| InternVL-2 | LIME Original | 0.063 0.030 | 6.864 4.421 |
@@
As shown in Tab 2, The MLLMs’ performances on the LIME are less than that on Original Bench on most tasks. Compared to the Origin benchmark, different MLLMs show a larger score range on our LIME, indicating that our LIME can better reflect the performance differences between models with a smaller amount of data.
Furthermore, we compare the score variations across different model series and model sizes. Figure 5 illustrates a clear positive correlation between model performance and model size within the same model series. Notably, LIME exhibits a more dispersed score distribution, effectively highlighting the differences in model performance. In Table 5.3 and 5.3, the Gini coefficient and standard deviation are used to measure the differences in overall score distribution across same model series and model sizes. The larger the Gini coefficient and standard deviation, the greater the disparity in data distribution. It can be observed that, whether within the same model series or the same model size, LIME achieves higher Gini and standard deviation values compared to the original bench. This indicates that LIME can better differentiate the performance differences between various models.
@@
According to the results in Figure 6 (left), LLMs tend to score higher on ScienceQA than on AI2D. This is because the former includes many commonsense text-only questions, such as ”What is the capital of Massachusetts?” which are unrelated to the images.
@@ 5.4 The Impact of Detail Image Perception
| LLaMA3-8B | 23.5 | 6.4 | 2.8 | 12.9 | 4.2 | 17.4 | 32.1 | 16.4 | 5.3 | 17.9 | |
| LLaMA3-70B | 24.0 | 7.7 | 3.3 | 12.3 | 4.3 | 21.8 | 38.1 | 39.4 | 6.0 | 22.0 | |
| Qwen1.5-32B | 28.8 | 6.7 | 6.5 | 9.4 | 4.0 | 4.7 | 39.3 | 46.6 | 9.2 | 13.7 | |
| Qwen1.5-72B | 25.4 | 2.5 | 3.2 | 10.1 | 4.1 | 7.8 | 42.7 | 44.2 | 6.0 | 15.2 | |
| Qwen2-7B | 27.6 | 6.7 | 6.9 | 11.2 | 4.1 | 15.0 | 44.2 | 45.5 | 12.5 | 19.0 | |
| Qwen2-72B | 26.3 | 6.9 | 2.7 | 10.8 | 4.4 | 10.6 | 36.3 | 45.2 | 5.2 | 16.8 | |
@@
In our data cleaning process, we removed many questions that most models could answer, as well as a small number of questions that are difficult for both humans and GPT-4 to answer, in order to make the benchmark better highlight the differences in model capabilities. As shown in Tab 5, to investigate whether the remaining samples need to be answered by using textual and image information, we conducted experiments using LLMs to generate answers on both the Original Benchmark and MLLMs Benchmark under QVD (question + visual description) setting.
@@ 6 Related work
In recent years, there has been increasing attention on establishing evaluation benchmarks to assess the performance of MLLMs in different scenarios to guide the development of MLLMs. Early multimodal evaluation benchmarks primarily focused on single tasks, such as Visual Question Answering (VQA) (Antol et al., 2015; Goyal et al., 2017; Kafle & Kanan, 2017; Singh et al., 2019; Marino et al., 2019), Image Captioning (Agrawal et al., 2019), and Information Retrieval (Wei et al., 2023). As MLLMs develop, simple benchmarks are no longer sufficient to evaluate the versatile capabilities of these models comprehensively, since most MLLMs have exceptional ability on those benchmarks. Consequently, numerous more difficult and diverse evaluation benchmarks have emerged in recent years to assess the capabilities of MLLMs comprehensively. For instance, MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024), are comprehensive benchmark tests for university-level multidisciplinary multimodal understanding and reasoning, MMBench (Liu et al., 2023c) has developed a comprehensive evaluation pipeline that offers fine-grained capability assessment and robust evaluation metrics. MMRA (Wu et al., 2024b) systematically establishes an association relation system among images to access the multi-image relation mining ability of MLLMs.
However, those benchmarks cannot distinguish the performance gaps among different models excellently, due to they still have some too simple or difficult samples that most models have the same results. Furthermore, training datasets across different models may contain the samples of those benchmarks, which results in data leakage issues (Fu et al., 2023). Mmstar (Chen et al., 2024a) and MMLU_Redux (Gema et al., 2024) have identified several issues within current benchmarks. Mmstar (Chen et al., 2024a) proposes an automated pipeline to filter benchmark data, aiming to detect potential data leakage, while MMLU_Redux (Gema et al., 2024) focuses on correcting annotation errors. However, there is still a pressing need for a comprehensive pipeline that fully addresses the challenges posed by multimodal datasets. In response to this, we introduce LIME: LESS IS MORE FOR MLLM EVALUATION. We have carefully selected six task types from existing mainstream benchmarks and scaled them down according to clear guidelines. This streamlined version retains the core elements of mainstream MLLM benchmarks, providing a more efficient and focused evaluation.
@@ 7 Conclusion
As MLLMs continue to advance, a notable absence of convenient and high-quality multimodal benchmarks has emerged. In response to this, we propose a pipeline aimed at semi-automatically refining existing benchmarks to enhance their quality, culminating in the development of LIME, which comprises 9,400 evaluation samples across 6 types of tasks and 10 different benchmark datasets. By refining the original benchmarks to filter question difficulty and eliminate potentially problematic items, LIME offers a more rigorous evaluation for MLLMs, necessitating a deeper understanding of image information. The outcomes of our evaluation experiments demonstrate the heightened challenge posed by LIME for MLLMs. We anticipate that our approach will contribute to the advancement of MLLM evaluation systems, and we are committed to continually enriching LIME with an expanded array of datasets through regular updates and expansions. Our ultimate goal is to provide the community with a simpler, more efficient, and more accurate evaluation method and suite for MLLMs.
@@ Appendix A Appendix
A.2 Pipeline details
A.2.1 prompt template details

### v3

@@ 5.2 Correlation Analysis
Figure 4 illustrates the correlation between the various sub-tasks in LIME and WildVision Bench. Most tasks in LIME exhibit a strong positive correlation with WildVision Bench. Six subtasks have correlation scores exceeding 80%. Additionally, the overall score of LIME correlates at 91% with WV-Elo, which is higher than any individual sub-task and the original bench’s correlations, demonstrating that the overall score of LIME provides a more comprehensive reflection of MLLMs’ capabilities.
@@
As an early foundational problem, the captioning task is extensively studied, and MLLMs demonstrate exceptional ability in this task. For instance, earlier models like InstructBlip perform exceptionally well on captioning tasks, and there is a broad presence of training data for image captioning in MLLMs’ training processes. However, the captioning task shows a negative correlation with all other sub-tasks. This indicates that previous metrics (e.g., BLEU, CIDEr) only focus on the overlap between the model-generated responses and the ground truth, but do not consider that MLLMs might generate content that is semantically close to the ground truth (i.e., the model-generated response may be semantically similar to the ground truth but expressed differently, or the model may generate more detailed content about the image). Consequently, we exclude it from the overall score calculation.
@@
On the one hand, the relevance of TextVQA, InfographicVQA, and OCRBench is relatively high. As shown in Fig. 4, the correlation of these tasks all surpasses 85%, and these two VQA tasks require MLLMs to understand fine-grained content in images to answer questions. This demonstrates that OCR tasks also rely on the ability of MLLMs to perceive fine-grained objective facts in images. On the other hand, POPE, ChartQA, and InfographicVQA all require reasoning abilities using extra commonsense knowledge. The correlation scores of these tasks are all above 70%, and POPE requires the model to use extra knowledge to solve the hallucination of MLLMs. We assume that ChartQA and infoVQA may also necessitate the use of additional common knowledge by the models to solve problems.
@@ 5.3 Effectiveness of LIME
| InternVL-2 | LIME Original | 0.061 0.030 | 6.972 4.421 |
@@
As shown in Tab 2, the MLLMs’ performances on LIME are less than those on the Original Bench for most tasks. Compared to the Origin benchmark, different MLLMs show a larger score range on our LIME, indicating that our LIME can better reflect the performance differences between models with a smaller amount of data.
Furthermore, we compare the score variations across different model series and model sizes. Figure 5 illustrates a clear positive correlation between model performance and model size within the same model series. Notably, LIME exhibits a more dispersed score distribution, effectively highlighting the differences in model performance. In Tab 5.3 and 5.3, the Gini coefficient and standard deviation are used to measure the differences in overall score distribution across the same model series and model sizes. The larger the Gini coefficient and standard deviation, the greater the disparity in data distribution. It can be observed that, whether within the same model series or the same model size, LIME achieves higher Gini and standard deviation values compared to the original bench. This indicates that LIME can better differentiate the performance differences between various models.
@@ 5.4 The Impact of Detail Image Perception
| LLaMA3-8B | 23.5 | 6.4 | 2.8 | 12.9 | 9.2 | 17.4 | 32.1 | 16.4 | 5.3 | 17.9 | |
| LLaMA3-70B | 24.0 | 7.7 | 3.3 | 12.3 | 9.3 | 21.8 | 38.1 | 39.4 | 6.0 | 22.0 | |
| Qwen1.5-32B | 28.8 | 6.7 | 6.5 | 9.4 | 8.7 | 4.7 | 39.3 | 46.6 | 9.2 | 13.7 | |
| Qwen1.5-72B | 25.4 | 2.5 | 3.2 | 10.1 | 8.9 | 7.8 | 42.7 | 44.2 | 6.0 | 15.2 | |
| Qwen2-7B | 27.6 | 6.7 | 6.9 | 11.2 | 8.9 | 15.0 | 44.2 | 45.5 | 12.5 | 19.0 | |
| Qwen2-72B | 26.3 | 6.9 | 2.7 | 10.8 | 9.6 | 10.6 | 36.3 | 45.2 | 5.2 | 16.8 | |
@@
In our data cleaning process, we remove many questions that most models can answer, as well as a small number of questions that are difficult for both humans and GPT-4V to answer, in order to make the benchmark better highlight the differences in model capabilities. As shown in Tab 5, to investigate whether the remaining samples need to be answered by using textual and image information, we conduct experiments using LLMs to generate answers on both the Original Benchmark and MLLMs Benchmark under QID (question + image description) setting.
@@
5.5 Existing benchmark still differs from real-world query.
To further investigate the gap between LIME and real-world users’ queries, we construct a similarity search system that compares them. MixEval (Ni et al., 2024) uses SentenceTransformers(Reimers, 2019) as the retrieval model, while Uniir (Wei et al., 2023) employs multimodal models like CLIP and BLIP. We use WildVision-Chat as the query data source, which contains 45.2k high-quality user questions, and employ SentenceTransformers to retrieve the top 10 most similar samples from LIME. To fully incorporate image information, we combine the question and image description as the query input. Additionally, we utilize Qwen2-72B to ensure a high level of relevance in the final results. As a result, we obtain a LIME-fit dataset containing 1.1k relevant samples. Existing benchmark can’t cover all types of real-world query.
In Figure 9, we compare the category distribution differences between LIME-fit and the WildVision Bench. It is evident that LIME-fit concentrates in a few specific categories (e.g., data analysis, general description, object recognition). However, it does not include instructions for solving real-world problems, such as Face Recognition, Problem Solving, and Scene Description. Furthermore, Figure 10 shows the frequency distribution of each subcategory in LIME-fit, which follows a long-tail distribution. This indicates that the current benchmark does not fully cover the instruction requirements of real-world scenarios.
@@ 6 Related work
In recent years, there has been increasing attention on establishing evaluation benchmarks to assess the performance of MLLMs in different scenarios to guide the development of MLLMs. Early multimodal evaluation benchmarks primarily focused on single tasks, such as Visual Question Answering (VQA)(Antol et al., 2015; Goyal et al., 2017; Kafle & Kanan, 2017; Singh et al., 2019; Marino et al., 2019), Image Captioning(Agrawal et al., 2019), and Information Retrieval (Wei et al., 2023). As MLLMs develop, simple benchmarks are no longer sufficient to evaluate the versatile capabilities of these models comprehensively, since most MLLMs demonstrate exceptional ability on those benchmarks. Consequently, numerous more difficult and diverse benchmarks have emerged in recent years to assess the capabilities of MLLMs comprehensively. For instance, MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024a) are comprehensive benchmark tests for university-level multidisciplinary multimodal understanding and reasoning. MMBench (Liu et al., 2023c) has developed a comprehensive evaluation pipeline that offers fine-grained capability assessment and robust evaluation metrics. MMRA (Wu et al., 2024b) systematically establishes an association relation system among images to assess the multi-image relation mining ability of MLLMs.
However, those benchmarks cannot distinguish the performance gaps among different models excellently, as they still contain some too simple or difficult samples that most models yield the same results on. Furthermore, training datasets across different models may contain the samples of those benchmarks, which results in data leakage issues (Fu et al., 2023). Mmstar (Chen et al., 2024a) and MMLU_Redux (Gema et al., 2024) have identified several issues within current benchmarks. Mmstar (Chen et al., 2024a) proposes an automated pipeline to filter benchmark data, aiming to detect potential data leakage, while MMLU_Redux (Gema et al., 2024) focuses on correcting annotation errors. However, there is still a pressing need for a comprehensive pipeline that fully addresses the challenges posed by multimodal datasets. In response to this, we introduce LIME: LESS IS MORE FOR MLLM EVALUATION. We have carefully selected six task types from existing mainstream benchmarks and scaled them down according to clear guidelines. This streamlined version retains the core elements of mainstream MLLM benchmarks, providing a more efficient and focused evaluation.
@@ 7 Conclusion
As MLLMs continue to advance, a notable absence of convenient and high-quality multimodal benchmarks has emerged. In response to this, we propose a pipeline aimed at semi-automatically refining existing benchmarks to enhance their quality, culminating in the development of LIME, which comprises 9,403 evaluation samples across 6 types of tasks and 10 different benchmark datasets. By refining the original benchmarks to filter question difficulty and eliminate potentially problematic items, LIME offers a more rigorous evaluation for MLLMs, necessitating a deeper understanding of image information. The outcomes of our evaluation experiments demonstrate the heightened challenge posed by LIME for MLLMs. We anticipate that our approach will contribute to the advancement of MLLM evaluation systems, and we are committed to continually enriching LIME with an expanded array of datasets through regular updates and expansions. Our ultimate goal is to provide the community with a simpler, more efficient, and more accurate evaluation method and suite for MLLMs.
@@ Appendix A Appendix
Figure 7shows the overall data distribution in LIME, and figure 8 shows an example for each category title
A.2 more experiment result
| T/F | Common VQA | InfoVQA | ScienceQA | OCR | Captioning | |||||||
| Model | Size | Overall | POPE | TVQA | OK VQA | ChQA | IVQA | AI2D | SciQA | OCRBen | COCO | TCaps |
| InternVL-2 2023 ( - ) | 40B | 66.85 | 51.69 | 77.98 | 19.45 | 88.33 | 73.92 | 69.20 | 98.33 | 75.87 | 63.10 | 48.94 |
| Qwen2-VL 2023b ( 1) | 7B | 65.28 | 53.05 | 74.56 | 28.18 | 83.17 | 78.50 | 58.20 | 75.00 | 77.61 | 68.74 | 110.60 |
| InternVL-1.5 2024b ( 1) | 26B | 64.12 | 51.69 | 69.88 | 35.47 | 87.00 | 70.92 | 54.81 | 91.67 | 60.65 | 69.24 | 112.63 |
| InternVL-2 2023 ( - ) | 26B | 63.98 | 54.63 | 75.20 | 16.08 | 87.67 | 70.58 | 62.80 | 96.33 | 71.09 | 76.18 | 56.91 |
| InternVL-2 2023 ( 1 ) | 8B | 62.00 | 49.21 | 66.10 | 24.20 | 83.75 | 68.25 | 60.40 | 95.67 | 70.65 | 42.55 | 25.44 |
| LLaVA-OneVision 2024 ( 1 ) | 7B | 61.95 | 52.37 | 65.22 | 37.32 | 80.83 | 68.17 | 59.20 | 92.00 | 47.83 | 104.74 | 108.18 |
| XComposer2-4KHD 2023 ( 4) | 7B | 57.52 | 46.28 | 60.30 | 28.13 | 80.42 | 66.17 | 54.90 | 70.00 | 53.04 | 97.07 | 78.07 |
| InternVL-2 2023( 1) | 4B | 57.22 | 47.18 | 62.29 | 17.48 | 81.92 | 60.50 | 54.00 | 94.33 | 67.17 | 35.99 | 21.67 |
| CogVLM-2 2024 ( 6) | 19B | 54.44 | 51.02 | 69.46 | 4.92 | 80.33 | 59.50 | 45.00 | 84.00 | 68.26 | 23.67 | 34.01 |
| Qwen2-VL 2023b ( 5) | 2B | 54.00 | 50.79 | 70.70 | 16.87 | 67.50 | 65.00 | 42.90 | 58.00 | 68.04 | 75.06 | 101.72 |
| InternVL-2 2023 ( 2) | 2B | 53.64 | 50.79 | 59.56 | 21.87 | 71.75 | 54.00 | 45.80 | 92.33 | 67.39 | 51.95 | 42.59 |
| CogVLM-1 2023 ( 1) | 17B | 51.03 | 55.10 | 71.20 | 31.70 | 61.67 | 57.25 | 31.40 | 53.67 | 41.96 | 29.28 | 38.56 |
| Cambrian 2024 ( 5) | 34B | 50.17 | 49.44 | 57.28 | 22.03 | 71.83 | 43.17 | 54.80 | 78.33 | 39.13 | 4.27 | 4.97 |
| Cambrian 2024 ( 4) | 13B | 48.57 | 50.79 | 58.93 | 24.13 | 69.25 | 42.83 | 45.20 | 62.67 | 42.39 | 7.40 | 6.52 |
| InternVL-2 2023 ( 3) | 1B | 48.21 | 52.82 | 57.55 | 15.37 | 65.83 | 46.25 | 37.00 | 84.33 | 65.00 | 15.19 | 13.19 |
| Cambrian 2024 ( 4) | 8B | 47.95 | 49.89 | 59.00 | 25.23 | 69.42 | 37.67 | 43.60 | 69.00 | 43.04 | 5.85 | 6.41 |
| LLaVA-1.6 2024 ( 3) | 34B | 44.06 | 47.00 | 51.20 | 10.40 | 64.33 | 42.08 | 49.60 | 64.67 | 37.17 | 84.25 | 48.25 |
| MiniCPM-LLaMA3-2.5 2024 ( 3) | 8B | 42.61 | 43.10 | 61.80 | 25.30 | 69.00 | 48.10 | 6.60 | 6.60 | 55.87 | 31.91 | 39.86 |
| LLaVA-OneVision 2024 ( 4) | 0.5B | 41.40 | 48.98 | 48.61 | 23.13 | 55.42 | 40.67 | 31.20 | 53.00 | 42.83 | 96.40 | 90.28 |
| LLaVA-LLaMA3 2023 ( 3) | 8B | 40.90 | 44.24 | 40.01 | 34.72 | 64.42 | 22.25 | 40.72 | 61.67 | 30.22 | 99.35 | 48.71 |
| Mantis-Idefics-2 2024 ( - ) | 8B | 39.25 | 44.24 | 44.51 | 29.07 | 59.00 | 20.50 | 35.80 | 70.00 | 32.17 | 61.82 | 103.07 |
| Deepseek-VL 2024 ( 2) | 7B | 38.10 | 48.50 | 44.80 | 25.00 | 54.67 | 22.33 | 37.00 | 68.33 | 25.43 | 54.22 | 83.21 |
| LLaVA-1.6-vicuna 2024 ( 4) | 13B | 37.08 | 43.10 | 43.90 | 16.10 | 54.50 | 28.75 | 38.10 | 53.00 | 31.96 | 76.62 | 47.83 |
| Idefics-2 2024 ( 2) | 8B | 36.39 | 42.00 | 56.50 | 35.60 | 13.08 | 23.92 | 38.10 | 78.67 | 42.61 | 61.23 | 94.51 |
| LLaVA-1.6-vicuna 2024 ( - ) | 7B | 30.15 | 41.10 | 39.00 | 12.50 | 43.08 | 22.67 | 27.10 | 47.33 | 23.70 | 76.05 | 48.35 |
| Mantis-SigLIP 2024 ( 1) | 8B | 29.13 | 45.60 | 26.34 | 32.43 | 35.33 | 16.25 | 27.70 | 62.67 | 10.65 | 68.16 | 81.21 |
| MiniCPM 2024 ( 2) | 1.0 | 26.15 | 44.00 | 37.00 | 6.20 | 35.75 | 13.42 | 27.90 | 60.67 | 14.57 | 68.96 | 76.65 |
| LLaVA-1.5 2023a ( 2) | 13B | 20.38 | 36.60 | 19.50 | 32.10 | 5.50 | 12.42 | 25.90 | 48.33 | 5.87 | 80.89 | 68.73 |
| LLaVA-1.5 2023a ( 1) | 7B | 17.20 | 32.51 | 16.50 | 23.43 | 5.25 | 9.08 | 24.05 | 49.00 | 4.78 | 79.20 | 65.73 |
| InstructBLIP-vicuna 2023 ( - ) | 7B | 15.55 | 45.10 | 11.40 | 22.10 | 3.00 | 9.08 | 21.90 | 34.33 | 4.35 | 102.08 | 53.14 |
| Tiny-LLaVA-1 2024 ( - ) | 1.4B | 13.95 | 37.00 | 18.70 | 0.90 | 4.50 | 12.17 | 22.80 | 44.67 | 3.48 | 63.19 | 58.91 |
| InternVL-2 | 40B | 80.31 | 89.23 | 82.59 | 50.98 | 85.52 | 76.08 | 85.88 | 98.56 | 79.90 | 99.15 | 62.03 |
| InternVL-1.5 | 26B | 79.49 | 88.90 | 79.00 | 60.70 | 83.70 | 72.50 | 78.90 | 94.50 | 71.40 | 95.80 | 148.10 |
| Qwen2-VL | 7B | 79.14 | 88.17 | 80.92 | 55.68 | 83.32 | 78.86 | 80.73 | 85.57 | 81.20 | 92.13 | 144.36 |
| InternVL-2 | 26B | 78.82 | 88.64 | 82.06 | 48.50 | 84.44 | 72.72 | 83.16 | 97.47 | 77.60 | 110.30 | 80.10 |
| LLaVA-OneVision | 7B | 78.71 | 89.17 | 76.02 | 60.98 | 80.12 | 70.69 | 81.38 | 95.88 | 62.10 | 140.45 | 136.97 |
| InternVL-2 | 8B | 77.84 | 87.90 | 77.00 | 52.02 | 82.48 | 70.65 | 82.25 | 97.03 | 76.50 | 89.77 | 36.70 |
| InternVL-2 | 4B | 73.97 | 87.71 | 74.51 | 38.43 | 81.04 | 65.19 | 78.08 | 96.03 | 75.00 | 54.08 | 30.17 |
| Cambrian | 34B | 73.26 | 88.46 | 72.11 | 52.07 | 74.60 | 51.48 | 80.41 | 85.52 | 59.00 | 8.18 | 6.08 |
| InternVL-2 | 2B | 73.00 | 88.90 | 72.39 | 43.74 | 74.72 | 57.69 | 72.70 | 94.25 | 75.50 | 79.52 | 59.81 |
| Cambrian | 13B | 72.39 | 88.53 | 73.07 | 53.28 | 72.60 | 50.73 | 73.93 | 79.08 | 61.40 | 14.33 | 9.44 |
| XComposer-4KHD | 7B | 71.93 | 87.00 | 74.30 | 51.90 | 80.60 | 72.80 | 34.40 | 96.00 | 66.90 | 134.00 | 111.40 |
| Cambrian | 8B | 71.84 | 88.24 | 72.47 | 52.17 | 73.44 | 48.05 | 72.99 | 80.32 | 61.60 | 9.13 | 7.97 |
| CogVLM-1 | 17B | 71.34 | 88.90 | 79.70 | 46.90 | 67.00 | 63.30 | 61.90 | 70.50 | 59.10 | 28.40 | 44.70 |
| MiniCPM-LLaMA3-2.5 | 8B | 71.22 | 88.00 | 75.00 | 52.30 | 72.90 | 56.90 | 71.70 | 53.00 | 69.40 | 35.50 | 52.90 |
| Qwen2-VL | 2B | 70.86 | 87.78 | 78.70 | 40.59 | 73.12 | 67.15 | 70.21 | 77.89 | 75.30 | 103.52 | 131.92 |
| CogVLM-2 | 19B | 69.93 | 87.56 | 77.59 | 18.51 | 79.84 | 62.62 | 72.41 | 90.93 | 76.60 | 24.10 | 42.23 |
| LLaVA-LLaMA3 | 8B | 69.74 | 87.80 | 65.40 | 60.20 | 69.30 | 37.60 | 71.60 | 73.30 | 55.00 | 135.00 | 69.60 |
| InternVL-2 | 1B | 68.46 | 87.94 | 69.67 | 33.84 | 71.40 | 52.02 | 62.56 | 89.59 | 74.20 | 49.34 | 18.03 |
| LLaVA-1.6-vicuna | 13B | 67.29 | 87.50 | 67.00 | 46.30 | 62.20 | 41.50 | 70.40 | 73.50 | 55.00 | 101.90 | 67.30 |
| LLaVA-1.6 | 34B | 67.22 | 85.60 | 68.90 | 31.00 | 67.40 | 51.90 | 76.10 | 82.70 | 58.60 | 114.40 | 69.10 |
| Mantis-Idefics-2 | 8B | 66.91 | 86.90 | 63.51 | 52.50 | 63.56 | 31.17 | 66.81 | 81.80 | 54.20 | 79.42 | 134.08 |
| Idefics-2 | 8B | 66.73 | 86.80 | 71.30 | 53.90 | 26.40 | 37.00 | 69.20 | 87.20 | 61.60 | 71.90 | 119.10 |
| LLaVA-OneVision | 0.5B | 65.65 | 88.33 | 65.85 | 44.17 | 61.36 | 46.23 | 57.09 | 67.03 | 57.60 | 131.90 | 120.81 |
| Deepseek-VL | 7B | 65.62 | 87.10 | 63.20 | 48.70 | 60.60 | 34.30 | 63.40 | 81.70 | 43.30 | 67.60 | 110.10 |
| LLaVA-1.6-vicuna | 7B | 64.80 | 87.60 | 64.90 | 44.20 | 55.00 | 37.00 | 65.30 | 70.20 | 52.40 | 100.00 | 72.00 |
| LLaVA-1.5 | 13B | 59.58 | 87.10 | 48.70 | 58.30 | 18.10 | 29.50 | 59.40 | 72.80 | 33.60 | 115.40 | 104.00 |
| Mantis-SigLIP | 8B | 58.96 | 81.47 | 49.59 | 52.90 | 42.56 | 26.56 | 57.84 | 75.36 | 34.50 | 91.37 | 111.43 |
| LLaVA-1.5 | 7B | 57.27 | 87.00 | 46.10 | 53.40 | 18.20 | 25.80 | 55.20 | 69.50 | 31.50 | 109.00 | 98.00 |
| MiniCPM | 1.0 | 56.18 | 85.10 | 55.30 | 47.30 | 15.40 | 20.10 | 56.90 | 43.00 | 60.00 | 25.90 | 41.60 |
| InstructBLIP-vicuna | 7B | 47.87 | 85.00 | 33.20 | 45.20 | 12.50 | 22.90 | 34.00 | 36.40 | 25.90 | 141.40 | 74.00 |
| Tiny-LLaVA-1 | 1.4B | 34.30 | 56.30 | 38.50 | 3.80 | 11.10 | 22.20 | 32.30 | 58.20 | 17.20 | 80.90 | 83.10 |
A.3 Pipeline details
A.3.1 prompt template details